# Implicit Neural Shape Optimization for 3D High-Contrast Electrical Impedance Tomography

### v1

@@ Abstract
We present a novel implicit neural shape optimization framework for 3D high-contrast Electrical Impedance Tomography (EIT), addressing scenarios where conductivity exhibits sharp discontinuities across material interfaces. These high-contrast cases, prevalent in metallic implant monitoring and industrial defect detection, challenge traditional reconstruction methods due to severe ill-posedness. Our approach synergizes shape optimization with implicit neural representations, introducing key innovations including a shape derivative-based optimization scheme that explicitly incorporates high-contrast interface conditions and an efficient latent space representation that reduces variable dimensionality. Through rigorous theoretical analysis of algorithm convergence and extensive numerical experiments, we demonstrate substantial performance improvements, establishing our framework as promising for practical applications in medical imaging with metallic implants and industrial non-destructive testing.
Keywords: High-Contrast Electrical Impedance Tomography, Shape Optimization, Shape Derivative, Implicit Neural Representation
@@ 1 Introduction
Electrical impedance tomography (EIT) has developed into a valuable imaging technique with promising applications in noninvasive medical diagnosis [17, 19] and non-destructive industrial testing [12, 23]. The technique aims to recover the conductivity distribution $\sigma\in L^{\infty}(D)$ from boundary measurements and is a severely ill-posed inverse problem. In this work, we focus on a specific formulation in which the conductivity exhibits sharp discontinuities across material interfaces, leading to the EIT interface reconstruction problem: recovery of a piecewise constant distribution characterized by an open subset $S\subset D$ with smooth boundary $\partial S$, expressed as $\sigma(x)=\sigma_{1}\chi_{D\setminus S}(x)+\sigma_{2}\chi_{S}(x)$. This interface reconstruction approach is particularly relevant in practical scenarios, including detection of metallic implants in biological tissues, localization of tumors with distinct conductivity properties, and identification of industrial defects - contexts where the primary interest lies in determining the geometric boundaries between regions of different electrical properties rather than reconstructing detailed conductivity maps.
A range of approaches have been developed to address the EIT interface reconstruction problem. Regularized Newton methods enhance stability through Tikhonov regularization [15], with extended approaches implementing edge-preserving total variation constraints [21]. Integral equation approaches [13] offer computational advantages by avoiding full-domain discretization while establishing Fréchet differentiability of the measurement operator. Shape optimization methods [14] directly evolve the interface geometry using gradient flows derived from shape calculus. The level-set method [30] represents a particularly powerful framework for interface evolution, offering natural handling of topological changes without prior knowledge of the solution structure.
@@
First proposed by Osher and Sethian for curve evolution, the level set methods gained prominence in EIT interface problems through pioneering work [36, 32, 26]. The level-set-based shape optimization method [7, 34] combines topological flexibility with rigorous mathematical frameworks, treating the interface $\Gamma=\partial S$ as a manifold with shape derivatives derived through adjoint analysis. This powerful approach unifies the mathematical rigor of shape calculus with the geometric adaptability of implicit representations, providing theoretical convergence guarantees through Fréchet differentiable shape derivatives [14] while elegantly handling topological changes. Computational advances include second-order curvature acceleration techniques [1] and edge-preserving total variation regularization [21], which significantly outperforms traditional Tikhonov methods in preserving sharp material interfaces. Recent extensions demonstrate the versatility of the method: multi-physics formulations [27], nonstationary approaches for resolving depth ambiguities [28], and incorporation of structural priors for complex industrial applications [25, 2]. This comprehensive framework, which forms the foundation for our present work, offers a robust solution for both clinical and industrial EIT applications.
Although level-set-based shape optimization methods have significantly advanced EIT interface reconstruction, critical limitations impede their application to comprehensive 3D problems. The explicit discretization of the interface inherent in conventional level set approaches [7] imposes prohibitive memory requirements, largely confining existing EIT studies to 2D or simplified quasi-3D geometries. Current heuristic regularization techniques [21] inadequately preserve complex anatomical features, frequently introducing geometric artifacts that compromise diagnostic precision. Moreover, these methods exhibit marked sensitivity to initial conditions and topological constraints during interface evolution, creating formidable convergence challenges. Recent breakthroughs in deep learning, particularly Implicit Neural Representations [31, 10], present promising pathways to overcome these obstacles. Neural implicit functions have exhibited exceptional capabilities in representing intricate 3D geometries through continuous, differentiable mappings [8]. Their capacity to encode compact latent representations while preserving fine geometric details makes them especially suitable for EIT interface reconstruction problems.
We develop the implicit neural shape optimization framework, which enhances traditional level-set-based shape optimization methods. The design of our algorithm is motivated by the need for an efficient and accurate representation of complex interfaces in EIT. By parameterizing the conductivity interface as $\Gamma=\{\bm{x}\in\mathbb{R}^{3}|f_{\theta}(\bm{x})=0\}$ using a differentiable neural network $f_{\theta}:\mathbb{R}^{3}\to\mathbb{R}$, we enable mesh-free optimization. By introducing a latent representation of surfaces, we significantly reduce the dimensionality of optimization variables, making the EIT problem more tractable with faster convergence and improved stability. We derive clear gradient expressions of the loss function with respect to these latent variables, enabling the design of efficient gradient descent algorithms. This integration of interface inverse problems with neural representation of surfaces not only advances 3D EIT reconstruction but also provides a robust framework for addressing general PDE-constrained shape optimization challenges. The key innovations and advantages of this work can be summarized as follows:
  • We introduce a novel implicit neural framework for EIT interface reconstruction that parameterizes complex conductivity interfaces using neural networks, effectively reducing the high-dimensional shape optimization problem to a more manageable latent space optimization.
  • We derive explicit gradient expressions of the objective function with respect to latent variables, enabling the design of an efficient gradient descent algorithm that navigates the latent space to identify optimal interface configurations.
  • We provide a rigorous mathematical analysis which establishes convergence guarantees for our optimization scheme, demonstrating that under appropriate conditions, the algorithm converges to stationary points of the regularized objective function.
@@ 2 The Problem Setting
$-\nabla\cdot(\sigma(x)\nabla u)=0\quad\text{in }\Omega,$ (1)
@@
We consider a high-contrast conductivity model [7], where conductivity exhibits sharp discontinuities across material interfaces. Let $D\subset\mathbb{R}^{3}$ be a simply connected bounded domain containing an unknown inclusion $S$ with a $C^{2}$-regular boundary, satisfying the geometric separation condition $\text{dist}(\partial S,\partial D)>\delta>0$ (see Figure 1). The high-contrast conductivity model is characterized by:
$\sigma(x)=\begin{cases}\sigma_{+}&x\in S\\ 1&x\in D\setminus\overline{S}\end{cases},\quad\text{with }\sigma_{+}\gg 1.$ (2)
@@
[…]
@@
The boundary measurements of the voltage $f=u|_{\Sigma}\in H^{1/2}(\Sigma)$ are then collected for the inverse problem.
@@ 2.1 Boundary Integral Formulation
This formulation reveals the severe ill-posedness inherent to the EIT problem. The off-diagonal blocks $\mathcal{S}_{\Sigma\Gamma}$ and $\mathcal{K}_{\Sigma\Gamma}$ that couple the inclusion boundary $\Gamma$ to the measurement boundary $\Sigma$ contain terms that decay exponentially with the distance between these boundaries, as illustrated in Figure 2. Specifically, the fundamental solution $G(\bm{x},\bm{y})$ decreases as $O(|\bm{x}-\bm{y}|^{-1})$, causing rapid attenuation of information from the interior interface to the exterior boundary. Consequently, small perturbations in boundary measurements can correspond to vastly different internal configurations, making the inverse problem extremely sensitive to measurement noise [4].
@@
3 Implicit Neural Shape Optimization
Our approach synergizes classical shape optimization theory with implicit neural representations, creating a powerful framework for reconstructing complex interfaces in EIT problems.
3.1 Implicit Neural Surfaces Representation and Topology Preservation
Traditional interface representation methods in EIT encounter significant challenges when applied to 3D reconstruction problems. Explicit mesh-based approaches, while intuitive, struggle with topological changes that naturally occur during reconstruction. Level-set methods offer greater flexibility but frequently introduce undesirable artifacts, particularly when working with noisy measurements, a common scenario in practical EIT applications.
@@ 3.1 Implicit Neural Surfaces Representation and Topology Preservation
While DeepSDF provides a powerful representation framework, they do not inherently guarantee topological consistency during shape manipulation or evolution. While DeepSDF provides a powerful representation framework, they do not inherently guarantee topological consistency during shape manipulation or evolution. This limitation is particularly problematic for EIT interface reconstruction, where preserving the correct topology of anatomical structures or industrial components is essential for accurate diagnosis or defect detection. To address this limitation, we adopt the Neural Diffeomorphic Flow (NDF) framework introduced by Sun et al. [37], which ensures robust topology preservation through carefully designed diffeomorphic constraints. The NDF approach conceptualizes shape deformation as a differentiable flow $\psi_{z}:\mathbb{R}^{3}\to\mathbb{R}^{3}$, parameterized by the latent code $z$. This flow transforms a template shape into the target shape while strictly preserving topological characteristics.
@@
For our EIT application, we leverage a pre-trained NDF model that has already learned to generate topologically consistent shapes within its latent space. This allows us to navigate the latent space during reconstruction while inheriting the topological guarantees provided by the diffeomorphic constraints — a critical advantage for accurate interface representation in both medical imaging and industrial inspection applications where topological errors could lead to misdiagnosis or false detections.
3.2 Optimization in Latent Space
By introducing the implicit neural representation of surfaces, we transform the challenging surface optimization problem on the infinite-dimensional admissible set $\mathcal{A}$ into a more tractable optimization problem in a compact latent space $\mathcal{M}$ (Figure 3). This reduction of dimensionality represents a paradigm shift in our approach to EIT interface reconstruction. Given a pre-trained NDF model, we seek the optimal latent code $z\in\mathcal{M}$ such that

### v2

# Generative Prior-Guided Neural Interface Reconstruction for 3D Electrical Impedance Tomography

@@ Abstract
Reconstructing complex 3D interfaces from indirect measurements remains a grand challenge in scientific computing, particularly for ill-posed inverse problems like Electrical Impedance Tomography (EIT). Traditional shape optimization struggles with topological changes and regularization tuning, while emerging deep learning approaches often compromise physical fidelity or require prohibitive amounts of paired training data. We present a transformative “solver-in-the-loop” framework that bridges this divide by coupling a pre-trained 3D generative prior with a rigorous boundary integral equation (BIE) solver. Unlike Physics-Informed Neural Networks (PINNs) that treat physics as soft constraints, our architecture enforces the governing elliptic PDE as a hard constraint at every optimization step, ensuring strict physical consistency. Simultaneously, we navigate a compact latent manifold of plausible geometries learned by a differentiable neural shape representation, effectively regularizing the ill-posed problem through data-driven priors rather than heuristic smoothing. By propagating adjoint shape derivatives directly through the neural decoder, we achieve fast, stable convergence with dramatically reduced degrees of freedom. Extensive experiments on 3D high-contrast EIT demonstrate that this principled hybrid approach yields superior geometric accuracy and data efficiency which is difficult to achieve using traditional methods, establishing a robust new paradigm for physics-constrained geometric discovery.
Keywords: Shape Optimization, Generative Prior, Electrical Impedance Tomography
@@ 1 Introduction
PDE-constrained shape optimization problems represent a fundamental class of inverse problems with broad applications across scientific computing, medical imaging, and engineering design. These problems seek to recover geometric interfaces or boundaries from indirect measurements, where the relationship between geometry and observations is governed by partial differential equations. The mathematical challenge lies in the severe ill-posedness of these inverse problems, compounded by the high-dimensional nature of shape spaces and the computational complexity of repeatedly solving forward PDEs during optimization. Electrical impedance tomography (EIT) exemplifies this challenge, with transformative applications in noninvasive medical diagnosis [17, 19] and non-destructive industrial testing [12, 24]. The technique aims to recover conductivity distributions from boundary measurements, with the high-contrast interface reconstruction problem being especially critical—where conductivity exhibits sharp discontinuities across material boundaries characterized by piecewise constant distributions $\sigma(x)=\sigma_{1}\chi_{D\setminus S}(x)+\sigma_{2}\chi_{S}(x)$ with open subset $S\subset D$ and smooth boundary $\partial S$. This formulation is particularly relevant for applications including metallic implant monitoring in biological tissues, tumor localization with distinct electrical properties, and industrial defect detection, where accurate geometric boundary determination is paramount for diagnostic and safety outcomes.
@@
A range of approaches have been developed to address the EIT interface reconstruction problem, with regularized Newton methods enhancing stability through Tikhonov regularization [15] and extended edge-preserving total variation constraints [22], while integral equation approaches [13] offer computational advantages by avoiding full-domain discretization and establishing Fréchet differentiability of the measurement operator. Shape optimization methods [14] directly evolve interface geometry using gradient flows derived from shape calculus, with the level-set method [34] representing a particularly powerful framework that gained prominence in EIT through pioneering work [40, 36, 28]. Level-set-based shape optimization [7, 38] combines topological flexibility with rigorous mathematical frameworks, treating the interface $\Gamma=\partial S$ as a manifold with shape derivatives derived through adjoint analysis, providing theoretical convergence guarantees through Fréchet differentiable shape derivatives while elegantly handling topological changes. Computational advances include second-order curvature acceleration techniques [1], edge-preserving total variation regularization that significantly outperforms traditional Tikhonov methods in preserving sharp material interfaces, and recent extensions demonstrating versatility through multi-physics formulations [29], nonstationary approaches for resolving depth ambiguities [30], and incorporation of structural priors for complex industrial applications [26, 2]. However, critical limitations impede their application to comprehensive 3D problems: explicit discretization imposes prohibitive memory requirements largely confining existing EIT studies to 2D or simplified quasi-3D geometries, current heuristic regularization techniques inadequately preserve complex anatomical features while frequently introducing geometric artifacts that compromise diagnostic precision, and these methods exhibit marked sensitivity to initial conditions and topological constraints during interface evolution, creating formidable convergence challenges.
Two dominant deep learning paradigms have emerged for addressing inverse problems, especially electrical impedance tomography (EIT) reconstruction problems. End-to-end learning approaches [20, 33] directly map boundary voltage measurements to conductivity distributions through deep neural networks, leveraging convolutional architectures, U-Net structures, and autoencoder frameworks to achieve remarkable computational efficiency compared to traditional iterative methods. However, these methods typically require vast quantities of paired training data (ground truth conductivity distributions and corresponding boundary measurements) and exhibit limited interpretability as ”black boxes.” Physics-Informed Neural Networks (PINNs) [31, 43] represent the alternative paradigm, parameterizing unknown conductivity fields with neural networks while incorporating the governing elliptic PDE as soft constraints through residual minimization in the loss function. While PINN extensions for EIT have shown promising results in handling complex geometries and multi-frequency measurements, training remains sensitive to the complex interplay between different loss terms (measurement misfit versus PDE residual) and hyperparameter tuning, often leading to convergence difficulties or physically inconsistent conductivity reconstructions. A complete introduction can be found in [11].
Recent advances in generative modeling have opened new avenues for solving inverse problems by learning data-driven priors from datasets. Generative approaches learn statistical regularities of feasible solutions and constrain optimization to learned manifolds of realistic configurations. However, existing generative approaches often struggle with incorporating complex physics constraints while maintaining computational efficiency. Our framework represents a principled hybrid architecture that occupies a distinct and highly advantageous middle ground between these two dominant paradigms by leveraging 3D generative models. Rather than learning explicit generative models of conductivity distributions, we employ neural implicit functions as learned geometric priors that inherently encode statistical regularities of anatomical shapes while maintaining differentiability for physics-based optimization. Neural implicit functions, which represent geometry as the level set of a continuous and differentiable mapping learned by a neural network, have shown exceptional capability in capturing intricate 3D shapes [8]. Unlike end-to-end methods, our approach is far more data-efficient, leveraging a pre-trained shape model but performing the actual inversion through a physics-based optimization loop, obviating the need for extensive paired measurement-shape data—a critical advantage in medical imaging where such data is scarce or impossible to acquire. In contrast to PINNs, our framework elegantly avoids the convergence pitfalls by maintaining the elliptic PDE governing the electric potential as a hard constraint, rigorously enforced at each optimization step by a dedicated, high-fidelity BIE solver, while the neural network’s role is judiciously confined to representing the geometric interface. This modular architecture represents more than a mere design choice; it is a philosophical advantage for scientific machine learning, enabling independent improvement and substitution of each component while ensuring every proposed solution is physically consistent. By parameterizing the conductivity interface as $\Gamma=\{\bm{x}\in\mathbb{R}^{3}|f_{\theta}(\bm{x})=0\}$ using a differentiable neural network $f_{\theta}:\mathbb{R}^{3}\to\mathbb{R}$, our framework achieves a fundamental paradigm shift by being inherently three-dimensional and completely mesh-free—eliminating the dimensional constraints that have historically limited EIT reconstruction to simplified geometries. This integration of interface inverse problems with neural representation of surfaces significantly reduces the dimensionality of optimization variables through latent representations, making the EIT problem more tractable with faster convergence and improved stability, while providing explicit gradient expressions that enable efficient gradient descent algorithms for navigating the latent space to identify optimal interface configurations.
Compared with classical level-set methods and Newton schemes with shape Hessians that optimize directly in mesh/level-set spaces with hand-tuned regularizers, our latent-space shape calculus brings some concrete advantages tailored to 3D EIT: It reduces the optimization from infinite geometric degrees of freedom to a compact latent vector, yielding substantial memory and time savings per BIE-based forward/adjoint solve; The 3D generative model produces smooth, well-defined surface normals almost everywhere, suppressing discretization artifacts that destabilize shape derivatives; The generative model’s diffeomorphic flow preserves topology during updates, avoiding the spurious splits/merges often observed in unconstrained level sets; We push adjoint shape-derivative integrands through the decoder to obtain latent-space gradients aligned with the learned manifold, improving conditioning and step efficiency. Empirically, these properties translate into marked stability and efficiency gains: we observe effective parameter reduction with rapid early descent and robust convergence under 20% noise, and consistently low Hausdorff and volume errors in the reconstruction of complex geometries(Figures 5, 7), while classical level-set pipelines typically require heavy TV/Tikhonov tuning and remain sensitive to initialization; In short, our approach keeps the PDE as a hard constraint via a matched BIE forward/adjoint pair but confines updates to a topology-preserving generative manifold, delivering robustness and data efficiency beyond traditional shape optimization.
The key innovations and advantages of this work can be summarized as follows:
  • Principled Hybrid Architecture: A novel ”solver-in-the-loop” framework that strategically combines pre-trained neural shape priors with physics-based optimization, achieving superior data efficiency without requiring extensive paired measurement-shape training data.
  • Hard Physics Constraint Enforcement: Maintains governing elliptic PDEs as hard constraints through dedicated BIE solvers at each optimization step, ensuring physical consistency and avoiding the convergence difficulties of soft constraint approaches.
  • Implicit Neural Interface Representation: Parameterizes conductivity interfaces as level sets of neural implicit functions, enabling inherently three-dimensional, completely mesh-free reconstruction that eliminates dimensional constraints and discretization artifacts.
  • Efficient Latent Space Optimization: Significantly reduces optimization dimensionality through neural latent representations, providing explicit gradients for efficient navigation of the solution space with improved convergence and stability.
@@ 2 The Problem Setting
[…]
@@
We consider a high-contrast conductivity model [7] that captures the essential physics of EIT interface reconstruction problems. In this formulation, the conductivity exhibits sharp discontinuities across material interfaces, creating the challenging inverse problem of boundary detection from electrical measurements.
Let $D\subset\mathbb{R}^{3}$ be a simply connected bounded domain containing an unknown inclusion $S$ with a $C^{2}$-regular boundary. To ensure well-posedness of the forward problem, we impose the geometric separation condition $\text{dist}(\partial S,\partial D)>\delta>0$ for some fixed $\delta>0$ (see Figure 1). The high-contrast conductivity distribution is then characterized by the piecewise constant function:
$\sigma(x)=\begin{cases}\sigma_{+}&x\in S\\ 1&x\in D\setminus\overline{S}\end{cases},\quad\text{with }\sigma_{+}\gg 1,$ (2)
where the large contrast ratio $\sigma_{+}\gg 1$ represents scenarios such as metallic implants in biological tissue or highly conductive inclusions in industrial materials.
The resulting conducting domain $\Omega=D\setminus\overline{S}$ exhibits a complex geometry: it is bounded externally by the accessible boundary $\Sigma:=\partial D$, where electrodes can be placed for current injection and voltage measurement, and internally by the unknown interface $\Gamma:=\partial S$ that we seek to reconstruct. We define the unit normal $n$ on the boundary $\partial\Omega=\Sigma\cup\Gamma$ to be the outward-pointing normal to the conducting domain $\Omega$. This geometric configuration encapsulates the fundamental challenge of EIT: inferring the internal structure $\Gamma$ from boundary observations on $\Sigma$, a severely ill-posed inverse problem that requires sophisticated regularization techniques for stable reconstruction.
@@
The forward problem in EIT involves solving a series of elliptic boundary value problems in the conducting domain $\Omega$. For our analysis, we focus on the current injection scenario, which corresponds to the typical experimental setup. Given current patterns $g\in H^{-1/2}(\Sigma)$ satisfying $\int_{\Sigma}g\,ds=0$ (current conservation), the governing equations for the electric potential $u$ are:
@@
As the contrast ratio approaches infinity, the electric field inside the highly conductive inclusion becomes negligible, making the potential $u$ approximately constant on the interface $\Gamma$. By fixing the gauge, this constant can be set to zero, which leads to the homogeneous Dirichlet boundary condition on $\Gamma$. The boundary measurements of the voltage $f=u|_{\Sigma}\in H^{1/2}(\Sigma)$ are then collected for the inverse problem.
@@ 2.1 Boundary Integral Formulation
This formulation reveals the severe ill-posedness inherent to the EIT problem. While the fundamental solution $G(\bm{x},\bm{y})$ decreases algebraically as $O(|\bm{x}-\bm{y}|^{-1})$, the off-diagonal blocks $\mathcal{S}_{\Sigma\Gamma}$ and $\mathcal{K}_{\Sigma\Gamma}$ map information between strictly disjoint boundaries $\Gamma$ and $\Sigma$. Because the integral kernels are infinitely smooth for $\bm{x}\neq\bm{y}$, these operators are highly compact [27]. Consequently, their singular values decay exponentially, which physically corresponds to the exponential attenuation of high-frequency spatial components of the interior information across the distance [4]. As a result, small perturbations in boundary measurements can correspond to vastly different internal configurations, making the inverse problem severely ill-posed and extremely sensitive to measurement noise. [4].
@@
3 Generative Prior-Guided Neural Interface Reconstruction
The EIT inverse problem is severely ill-posed, requiring regularization to obtain unique and stable solutions. Rather than relying on hand-crafted priors (e.g., Tikhonov or Total Variation), we employ a pre-trained 3D generative model as a learned prior over plausible interfaces. This shifts the problem from unconstrained evolution in an infinite-dimensional space to a principled search on a compact, data-driven shape manifold. By integrating classical shape optimization with generative priors, we establish a rigorous framework for PDE-constrained interface reconstruction in 3D.
3.1 3D Generative Models as Learned Shape Priors
A critical component of effective EIT reconstruction is the choice of prior over interfaces. Traditional 3D approaches face notable difficulties under noise and sparse measurements. Mesh-based methods struggle with evolving complex geometries, and level-set methods can introduce artifacts. In contrast, a pre-trained 3D generative model provides a data-driven prior that concentrates the search on plausible shapes.
@@ 3 Generative Prior-Guided Neural Interface Reconstruction
This latent code $z$ becomes the optimization variable in our EIT reconstruction, allowing us to work in a compact representation space rather than directly manipulating high-dimensional geometric parameters.
While DeepSDF provides a powerful representation framework, it does not inherently guarantee topological consistency during shape manipulation or evolution. This limitation is particularly problematic for EIT interface reconstruction, where preserving the correct topology of anatomical structures or industrial components is essential for accurate diagnosis or defect detection. To address this limitation, we adopt the Neural Diffeomorphic Flow (NDF) framework introduced by Sun et al. [41], which ensures robust topology preservation through carefully designed diffeomorphic constraints. The NDF approach conceptualizes shape deformation as a differentiable flow $\psi_{z}:\mathbb{R}^{3}\to\mathbb{R}^{3}$, parameterized by the latent code $z$. This flow transforms a template shape into the target shape while strictly preserving topological characteristics.
@@
For our EIT application, we employ a pre-trained generative model that has already learned to generate topologically consistent shapes within its latent space, providing us with a learned shape prior. This allows us to navigate the latent space during reconstruction while inheriting the topological guarantees provided by the diffeomorphic constraints — a critical advantage for accurate interface representation in both medical imaging and industrial inspection applications where topological errors could lead to misdiagnosis or false detections.
3.2 Generative Prior-Guided Optimization on Learned Shape Manifolds
To address the severe ill-posedness, we transform the challenging shape optimization problem from the infinite-dimensional admissible set $\mathcal{A}$ into a principled search within the learned manifold of plausible shapes encoded in latent space $\mathcal{M}$ (Figure 3). This dimensionality reduction represents a paradigm shift in our approach to EIT interface reconstruction. Given a pre-trained 3D generative model, we seek the optimal latent code $z\in\mathcal{M}$ such that